OpenStudy (anonymous):

-6+m over 9< -5 suppose to be a line under this <

OpenStudy (anonymous):

\[\frac{m}{9}\leq -5+6\] \[\frac{m}{9}\leq 1\] \[m\leq 1\times 9\] \[m\leq 9\]

OpenStudy (anonymous):

thank you